ok, i have a lot of math that i don't understand at all. in class, i try to listen and understand it, but it's just too difficult for me. Please help!
i have 6 problems altogether. Here's 3:
1) 5x+7=3x-5
2)7+3x=5x-5
3)2.5x-8=5x+12
THANK YOU SOO MUCH!!
-brianne..idaho. 8th grade

Answer by stanbon(75887)

1) 5x+7=3x-5
Subtract 3x from both sides to get:
Subtract 7 from both sides to get:
2x=-12
x=-6
-----------------
2)7+3x=5x-5
Subtract 3x from each side:
Add 5 to both sides:
2x=12
x=6
---------------------------
3)2.5x-8=5x+12
Subtract 2.5x from both sides:
Subtract 12 from both sides:
2.5x=-20
Divide by 2.5 to get:
x=-20
Cheers,
Stan H.